# Common Data Format (CDF)—a Standardized Format for Match-Data in Football (Soccer)

### v2

@@ Abstract
During football matches, a variety of different parties (e.g., companies) each collect (possibly overlapping) data about the match ranging from basic information (e.g., starting players) to detailed positional data. This data is provided to clubs, federations, and other organizations who are increasingly interested in leveraging this data to inform their decision making. Unfortunately, analyzing such data pose significant barriers because each provider may (1) collect different data, (2) use different specifications even within the same category of data, (3) represent the data differently, and (4) delivers the data in a different manner (e.g., file format, protocol). Consequently, working with these data requires a significant investment of time and money. The goal of this work is to propose a uniform and standardized format for football data called the Common Data Format (CDF). The CDF specifies a minimal schema for five types of match data: match sheet data, video footage, event data, tracking data, and match meta data. It aims to ensure that the provided data is clear, sufficiently contextualized (e.g., its provenance is clear), and complete such that it enables common downstream analysis tasks. Concretely, this paper will detail the technical specifications of the CDF, the representational choices that were made to help ensure the clarity of the provided data, and a concrete approach for delivering data in CDF.
@@ 5 The Common Football Data Format
In essence, the CDF aims to present a precise and well-defined interface to football data that mitigates the existing pain-points associated with the current data landscape. That is, it should enable end-users such as clubs, federations and researchers, to easily build their own logic and applications on top of provided data without having to contend with some of the (arbitrary) design choices made by data collectors. Having vendors supply data in the CDF will help shorten the time between data exploration and delivery of actionable products to stakeholders as well as facilitate more impactful collaborations with researchers and developers.
This section describes the key properties of the CDF and discusses some of the design choices that were made. First, we will outline the technical specifications of the CDF in terms of what data sources are included and within each data source what information should be provided. Second, we will discuss some of the representational choices that were made to help ensure the clarity of the provided data. Third, we will propose a concrete way to deliver data in the CDF to enable easy ingestion and analysis of the data. Finally, we would like to highlight that this is the first version of the CDF and it will continue to evolve as the data landscape changes.
@@ 5.1 Specification of the CDF Schema
Table 4 shows the information that is required for basic center of body tracking. Beyond the player coordinates, each frame of the tracking data should be augmented with information about which team has the ball and whether the ball is in or out of play. Every period starts with $\text{\emph{frame}}=0$ where period can take on the values first half, second half, first half of overtime, second half of overtime, and penalty shootouts. If the ball is tracked at a different sampling rate, then the provider should synchronize this to the player sampling rate and only provide these measurements in the schema outlined in Table 4. In this situation, the provider can also provide an additional, separate file that omits player positions and just reports the ball location at its (higher) sampling rate. If skeletal tracking is provided, a separate schema shown in Table 5 is used for the joint locations. A separate schema is used for ease of processing.
@@
Root Field Description Type match id Unique match identifier String recording fps Frames per second (i.e. frame rate) from vendor Integer recording resolution Resolution of the video in pixels (e.g., 3840x2160, 1920x1080, …) String recording start_time The start time in UTC of the recording String recording operation_type Information whether the camera operation was manual or automated String recording recording_type Information how the video was recorded (e.g. fixed camera, camcorder,…) String recording perspective Camera angle (tactical_wide, camera_1, high_behind_right, high_behind_left, cable_camera, 16m_right, 16m_left, broadcast) String recording/camera x x location of the camera in relation to the pitch center (m) Float recording/camera y y location of the camera in relation to the pitch center (m) Float recording/camera z z location of the camera in relation to the pitch center (m) Float whistles/{i} type Whistles that start and end major periods of play such as the start and end of halves and interruptions (e.g., weather, VAR review, player health events, streakers or abandoned). Examples of types first_half, second_half, weather_delay, health_delay, injury_treatment fan_health_delay etc. String whistles/{i} sub_type Sub type related to an interruption, for example start or end. String whistles/{i} time The time in UTC of the whistle String whistles/{i} video_time The time tag of the whistle in milliseconds Integer
@@
Root Field Description Type match id Unique match identifier String meta is_synced Indicates if synced tracking data is available (true) or not (false) Boolean event id Unique identifier of the event String event time Absolute time in UTC of when the event started (e.g., moment the pass is given, the moment the ball leaves the hands on a thrown in) String event period Period of the match which can be first_half, second_half, first_half_extra, second_half_extra, or shootout String event type Name of the event type (e.g. shot, pass, referee, misc, etc.) String event sub_type Name of the event subtype, which can be for shot - (None, penalty_kick, free_kick, corner_kick); pass - (None, throw_in, free_kick, corner_kick, goal_kick, kick_off); referee - (final_whistle, foul, caution, offside, substitution, player_on, player_off. Player on and player off events occur for example when a player has to abandon the game due to a lack of available substitutions, or as a temporary measure after a medical check); misc - (other_ball_action, chance_without_shot, tackle) String event outcome Denotes whether the event was successful (true) or not (false) Boolean event outcome_detailed Detailed event outcome options: shot - (successful, saved, blocked, wide, woodwork, own_goal); pass - (e.g. successful, out_of_play, intercepted); referee - (start, end, injury, yellow_card, red_card or second_yellow_card); misc - (e.g. successful, unsuccessful) String event player_id Unique identifier of the player performing the action. For example, player committing a pass, or making a tackle. String event team_id Unique team identifier of the player performing the action String event receiver_id Unique identifier of the player receiving a pass. Leave null when the event is not a pass. String event receiver_time Absolute time in UTC of the first moment the ball was received. Leave null when the event is not a pass. String event x x location where the action of player_id happened (m). Float event y y location where the action of player_id happened (m). Float event x_end x location where the action of player_id ended (m). Float event y_end y location where the action of player_id ended (m). Float event body_part Denotes the body part used by player_id which can take on the values left_foot, right_foot, head, or other String event related_event_ids Unique identifier(s) of the events related to the action, or pass and associated receival event. For example, a related aerial duel event, the receival event related to a pass, or the player_off event associated with a player_on event. Leave null if no related events exist. Array[String]
@@
Root Field Description Type frame_id Unique frame identifier Integer period Period of the match (first_half, second_half, first_half_extratime, second_half_extratime, shootout) String match id Unique match identifier String teams/{home|away} id Unique identifier for the home or away team String teams/{home|away}/players/{i} id Unique identifier for a player String teams/{home|away}/players/{i} x x location of the player on the pitch (m). Float teams/{home|away}/players/{i} y y location of the player on the pitch (m). Float ball x x location of the ball on the pitch (m). Float ball y y location of the ball on the pitch (m). Float ball z z location of the ball on the pitch (m). Float
@@
Root Field Description Type frame_id Unique frame identifier Integer period Period of the match (first_half, second_half, first_half_extra, second_half_extra, shootout) String match id Unique match identifier String teams/{home|away} id Unique identifier for the home or away team String teams/{home|away}/players/{i} id Unique identifier for a player String teams/{home|away}/players/{i}/landmarks/{j} index Unique identifier for a landmark String teams/{home|away}/players/{i}/landmarks/{j} name Unique name for a landmark String teams/{home|away}/players/{i}/landmarks/{j} x Relative x coordinate of landmark in relation to the point of origin (m) Float teams/{home|away}/players/{i}/landmarks/{j} y Relative y coordinate of landmark in relation to the point of origin (m) Float teams/{home|away}/players/{i}/landmarks/{j} z Relative z coordinate of landmark in relation to the point of origin (m) Float teams/{home|away}/players/{i}/landmarks/{j} children List of children indexes associated with keypoint Array[Int] teams/{home|away}/players/{i}/landmarks/{j} is_visible If landmark is immediately visible or inferred Integer
@@
Root Field Description Type competition id Unique identifier for the competition String season id Unique identifier for the season String match id Unique identifier for the match String match kickoff_time Scheduled kickoff time in UTC String match/periods/{i} type Possible options are first_half, second_half, first_half_extratime, second_half_extratime, shootout String match/periods/{i} play_direction The direction of play for the home team. Possible options are left_right or right_left. String match/whistles/{i} type Whistles that start and end major periods of play such as the start and end of halves and interruptions (e.g., weather, VAR review, player health events, streakers or abandoned). Examples of types first_half, second_half, weather_delay, health_delay, injury_treatment fan_health_delay etc. String match/whistles/{i} sub_type Sub type related to an interruption, for example start or end. String match/whistles/{i} time The time in UTC of the whistle String teams/{home|away} id Unique identifier for the home or away team String teams/{home|away}/players/{i} Array of player objects. One array under teams/home, one array under teams/away Array teams/{home|away}/players/{i} id Unique player identifier String teams/{home|away}/players/{i} team_id Unique team identifier denoting the team the player plays for String teams/{home|away}/players/{i} jersey_number Jersey number for a player Integer teams/{home|away}/players/{i} is_starter Denotes whether a player started the game (true) or not (false) Boolean stadium id Unique identifier for the stadium String stadium pitch_length Length of the pitch (m), unknown if not available, null if not available Float stadium pitch_width Width of the pitch (m), unknown if not available, null if not available Float meta/video perspective Camera perspective (e.g. in_stadium, broadcast, tactical, tactical_wide). String meta/limb nodes Skeletal hierarchy used (i.e., what landmarks are tracked). A explicit example of how to provide this data can be found in Figure 6 JSON Object meta/{tracking|video|limb|ball|meta|cdf} version Version number for the data collection in use (e.g. ”0.1.0”) String meta/{tracking|video|limb|ball|meta} name Vendor name of the tracking, limb tracking, event or video data String meta/{tracking|video|limb|ball} fps Frames per second (i.e., frame rate) of tracking, limb tracking or video Integer meta/{tracking|event|limb|ball} collection_timing Indicates if the data was collected live or post_match. String
@@ 5.3 Structure of Delivered Data
The CDF covers data that are provided both post-match and real-time (i.e., live). Therefore the CDF will use a file format that supports both use cases in order to avoid needing separate processes and file structures for handling real-time data and post-match data. Concretely, data that is not streamed (e.g., match sheet data, meta data) should be delivered in the JSON format (.json) and real-time data (e.g., event data, tracking data) should delivered using JSON Lines (.jsonl). In all cases, data should always be encoded in UTF-8.
@@ 5.3.1 Meta- and Match Sheet Data File Format
Finally, if provided, skeletal tracking data should be accompanied by a skeletal hierarchy in the meta data under ”meta/limb_nodes”. It should adhere to the format shown in Figure 6 which shows a simplified hierarchy of a standardized T-pose. The format is borrowed from the glTF2.0 file format  [gltf2_specification]. The ”name” field specifies the joint under consideration, it follows a snake_case formatting with the ”right”, ”left” indicators as a suffix. The ”rotation” field is a quaternion defining the rotation of the joint in a T-pose, this can generally be left as [0.0, 0.0, 0.0, 1.0], but should be adjusted for joints that do not make right angles with their parent node (e.g. ”thumb_right” or ”eye_left”). Translation should follow the ”Y-up” approach such that $x$, $y$ and $z$ components specify the orientation in relation to their parent node. For example ”hip” has a child-node ”spine” (as defined by the landmark node indices in ”children”) and there for [0, 1, 0] indicates the spine goes up (1) in the Y-axis and does not deviate in the $x$ or $z$ direction compared to the hip.
@@ 5.3.2 Event and Tracking Data File Format
Figure 4 illustrates the first two events of the first half. Figures 5 shows a simplified example for two frames of player center of body tracking data. Note that we are only showing the mandatory CDF fields. These can be augmented with the optional fields (see Tables 8 and 9). If skeletal (limb) tracking is provided, it should follow the same structure as the tracking data JSON Lines format. Instead of a players $x$ and $y$ coordinates, the provider would provide coordinates for each landmark (e.g., $left\_hip\_x$, $right\_hip\_y$, $right\_hip\_x$) as specified by Table 5. The coordinate system should be the same used in the center of body tracking data, e.g., $x$, $y$, $z$ and are relative to the pitch. The field $is\_visible$ indicates if the part was actually measured, or inferred via an algorithm due to being hidden. Providers have flexibility over which body parts they track. This should be communicated beforehand.
@@ B.1 Optional Match Sheet Data
B.2 Optional Event Data
@@
B.3 Optional Tracking Data
@@
B.4 Optional Meta-Data

### v3

# The Common Data Format (CDF): A Standardized Format for Match-Data in Football (Soccer)

@@ Abstract
During football matches, a variety of different parties (e.g., companies) each collect (possibly overlapping) data about the match ranging from basic information (e.g., starting players) to detailed positional data. This data is provided to clubs, federations, and other organizations who are increasingly interested in leveraging this data to inform their decision making. Unfortunately, analyzing such data pose significant barriers because each provider may (1) collect different data, (2) use different specifications even within the same category of data, (3) represent the data differently, and (4) delivers the data in a different manner (e.g., file format, protocol). Consequently, working with these data requires a significant investment of time and money. The goal of this work is to propose a uniform and standardized format for football data called the Common Data Format (CDF). The CDF specifies a minimal schema for five types of match data: match sheet data, video footage, event data, tracking data, and match meta data. It aims to ensure that the provided data is clear, sufficiently contextualized (e.g., its provenance is clear), and complete such that it enables common downstream analysis tasks. Concretely, this paper will detail the technical specifications of the CDF, the representational choices that were made to help ensure the clarity of the provided data, and a concrete approach for delivering data in CDF. This represents version 1.0.0 of the CDF.
@@ 5 The Common Football Data Format
In essence, the CDF aims to present a precise and well-defined interface to football data that mitigates the existing pain-points associated with the current data landscape. Specifically, the CDF targets clubs, federations and researchers. It aims to allow them to easily build their own logic and applications on top of provided data without having to contend with some of the (arbitrary) design choices made by data collectors. Having vendors supply data in the CDF will help shorten the time between data exploration and delivery of actionable products to stakeholders as well as facilitate more impactful collaborations with researchers and developers.
This section presents Version 1.0.0 of the CDF. It describes the key properties of the CDF and discusses some of the design choices that were made. First, we will outline the technical specifications of the CDF in terms of what data sources are included and within each data source what information should be provided. Second, we will discuss some of the representational choices that were made to help ensure the clarity of the provided data. Third, we will propose a concrete way to deliver data in the CDF to enable easy ingestion and analysis of the data. Finally, we would like to highlight that this is the first version of the CDF and it will continue to evolve as the data landscape changes.
@@ 5.1 Specification of the CDF Schema
Table 4 shows the information that is required for basic center of body tracking. Beyond the player coordinates, each frame of the tracking data should be augmented with information about which team has the ball and whether the ball is in or out of play. Every period can take on the values first half, second half, first half of overtime, second half of overtime, and penalty shootouts. Frame identifiers should be monotonically increasing unique integers starting at 0. If the ball is tracked at a different sampling rate, then the provider should synchronize this to the player sampling rate and only provide these measurements in the schema outlined in Table 4. In this situation, the provider can also provide an additional, separate file that omits player positions and just reports the ball location at its (higher) sampling rate. If skeletal tracking is provided, a separate schema shown in Table 5 is used for the joint locations. A separate schema is used for ease of processing.
@@
Root Field Description Type match id Unique match identifier String recording fps Frames per second (i.e. frame rate) from vendor Integer recording resolution Resolution of the video in pixels (e.g., 3840x2160, 1920x1080, …) String recording start_time The start time in UTC of the recording String recording operation_type Information whether the camera operation was manual or automated String recording recording_type Information how the video was recorded (e.g. fixed camera, camcorder,…) String recording perspective Camera angle (tactical_wide, camera_1, high_behind_right, high_behind_left, cable_camera, 16m_right, 16m_left, broadcast) String whistles/{i} type Whistles that start and end major periods of play such as the start and end of halves and interruptions (e.g., weather, VAR review, player health events, streakers or abandoned). Examples of types first_half, second_half, weather_delay, health_delay, injury_treatment fan_health_delay etc. String whistles/{i} sub_type Sub type related to an interruption, for example start or end. String whistles/{i} time The time in UTC of the whistle String whistles/{i} video_time The time tag of the whistle in milliseconds Integer
@@
Root Field Description Type match id Unique match identifier String meta is_synced Indicates if synced tracking data is available (true) or not (false) for this event Boolean event id Unique identifier of the event String event time Absolute time in UTC of when the event started (e.g., moment the pass is given, the moment the ball leaves the hands on a thrown in) String event period Period of the match which can be first_half, second_half, first_half_extra, second_half_extra, or shootout String event type Name of the event type (e.g. shot, pass, referee, misc, etc.) String event sub_type Name of the event subtype, which can be for shot - (None, penalty_kick, free_kick, corner_kick); pass - (None, throw_in, free_kick, corner_kick, goal_kick, kick_off); referee - (final_whistle, foul, caution, offside, substitution, player_on, player_off. Player on and player off events occur for example when a player has to abandon the game due to a lack of available substitutions, or as a temporary measure after a medical check); misc - (other_ball_action, chance_without_shot, tackle) String event is_successful Denotes whether the event was successful (true) or not (false) Boolean event outcome_type Detailed event outcome options: shot - (successful, saved, blocked, wide, woodwork, own_goal); pass - (e.g. successful, out_of_play, intercepted); referee - (start, end, injury, yellow_card, red_card or second_yellow_card); misc - (e.g. successful, unsuccessful) String event player_id Unique identifier of the player performing the action. For example, player committing a pass, or making a tackle. String event team_id Unique team identifier of the player performing the action String event receiver_id Unique identifier of the player receiving a pass. Leave null when the event is not a pass or when the pass has no receiver. String event receiver_time Absolute time in UTC of the first moment the ball was received. Leave null when the event is not a pass or when the pass has no receiver. String event x x location where the action of player_id happened (m). Float event y y location where the action of player_id happened (m). Float event x_end x location where the action of player_id ended (m). Float event y_end y location where the action of player_id ended (m). Float event body_part Denotes the body part used by player_id which can take on the values left_foot, right_foot, head, or other String event related_event_ids Unique identifier(s) of the events related to the action, or pass and associated receival event. For example, a related aerial duel event, the receival event related to a pass, or the player_off event associated with a player_on event. Leave null if no related events exist. Array[String]
@@
Root Field Description Type frame_id Unique frame identifier Integer timestamp Timestamp of the frame in UTC String period Period of the match (first_half, second_half, first_half_extratime, second_half_extratime, shootout) String match id Unique match identifier String teams/{home|away} id Unique identifier for the home or away team String teams/{home|away}/players/{i} id Unique identifier for a player String teams/{home|away}/players/{i} x x location of the player on the pitch (m). Float teams/{home|away}/players/{i} y y location of the player on the pitch (m). Float ball x x location of the ball on the pitch (m). Float ball y y location of the ball on the pitch (m). Float ball z z location of the ball on the pitch (m). Float
@@
Root Field Description Type frame_id Unique frame identifier Integer timestamp Timestamp of the frame in UTC String period Period of the match (first_half, second_half, first_half_extra, second_half_extra, shootout) String match id Unique match identifier String teams/{home|away} id Unique identifier for the home or away team String teams/{home|away}/players/{i} id Unique identifier for a player String teams/{home|away}/players/{i}/landmarks/{j} index Unique identifier for a landmark String teams/{home|away}/players/{i}/landmarks/{j} name Name for a landmark String teams/{home|away}/players/{i}/landmarks/{j} x Relative x coordinate of landmark in relation to the point of origin (m) Float teams/{home|away}/players/{i}/landmarks/{j} y Relative y coordinate of landmark in relation to the point of origin (m) Float teams/{home|away}/players/{i}/landmarks/{j} z Relative z coordinate of landmark in relation to the point of origin (m) Float teams/{home|away}/players/{i}/landmarks/{j} children List of children indexes associated with keypoint Array[Int] teams/{home|away}/players/{i}/landmarks/{j} is_visible If landmark is detected (true) or inferred (false) Boolean
@@
Root Field Description Type competition id Unique identifier for the competition String season id Unique identifier for the season String match id Unique identifier for the match String match kickoff_time Scheduled kickoff time in UTC String match/periods/{i} type Possible options are first_half, second_half, first_half_extratime, second_half_extratime, shootout String match/periods/{i} play_direction The direction of play for the home team. Possible options are left_right or right_left. String match/whistles/{i} type Whistles that start and end major periods of play such as the start and end of halves and interruptions (e.g., weather, VAR review, player health events, streakers or abandoned). Examples of types first_half, second_half, weather_delay, health_delay, injury_treatment fan_health_delay etc. String match/whistles/{i} sub_type Sub type related to an interruption, for example start or end. String match/whistles/{i} time The time in UTC of the whistle String teams/{home|away} id Unique identifier for the home or away team String teams/{home|away}/players/{i} Array of player objects. One array under teams/home, one array under teams/away Array teams/{home|away}/players/{i} id Unique player identifier String teams/{home|away}/players/{i} team_id Unique team identifier denoting the team the player plays for String teams/{home|away}/players/{i} jersey_number Jersey number for a player Integer teams/{home|away}/players/{i} is_starter Denotes whether a player started the game (true) or not (false) Boolean stadium id Unique identifier for the stadium String stadium pitch_length Length of the pitch (m), unknown if not available, null if not available Float stadium pitch_width Width of the pitch (m), unknown if not available, null if not available Float meta/video perspective Camera perspective (e.g. in_stadium, broadcast, tactical, tactical_wide). String meta/landmarks nodes Skeletal hierarchy used (i.e., what landmarks are tracked). A explicit example of how to provide this data can be found in Figure 6 JSON Object meta/{event|tracking|video|landmarks|ball|meta|cdf} version Version number for the data collection in use (e.g. ”0.1.0”) String meta/{event|tracking|video|landmarks|ball|meta} name Vendor name of the tracking, limb tracking, event or video data String meta/{tracking|video|landmarks|ball} fps Frames per second (i.e., frame rate) of tracking, limb tracking or video Integer meta/{event|tracking|landmarks|ball} collection_timing Indicates if the data was collected live or post_match. String
@@ 5.3 Structure of Delivered Data
The CDF covers data that are provided both post-match and real-time (i.e., live). Therefore the CDF will use a file format that supports both use cases in order to avoid needing separate processes and file structures for handling real-time data and post-match data. Concretely, data that is not streamed (e.g., match sheet data, meta data) should be delivered in the JSON format (.json) and real-time data (e.g., event data, tracking data) should delivered using JSON Lines (.jsonl). [^13] We acknowledge that other end users such as the media may have other requirements with regards to live data delivery. In all cases, data should always be encoded in UTF-8.
@@ 5.3.1 Meta- and Match Sheet Data File Format
Finally, if provided, skeletal tracking data should be accompanied by a skeletal hierarchy in the meta data under ”meta/landmarks_nodes”. It should adhere to the format shown in Figure 6 which shows a simplified hierarchy of a standardized T-pose. The format is borrowed from the glTF2.0 file format  [gltf2_specification]. The ”name” field specifies the joint under consideration, it follows a snake_case formatting with the ”right”, ”left” indicators as a suffix. The ”rotation” field is a quaternion defining the rotation of the joint in a T-pose, this can generally be left as [0.0, 0.0, 0.0, 1.0], but should be adjusted for joints that do not make right angles with their parent node (e.g. ”thumb_right” or ”eye_left”). Translation should follow the ”Y-up” approach such that $x$, $y$ and $z$ components specify the orientation in relation to their parent node. For example ”hip” has a child-node ”spine” (as defined by the landmark node indices in ”children”) and there for [0, 1, 0] indicates the spine goes up (1) in the Y-axis and does not deviate in the $x$ or $z$ direction compared to the hip.
@@ 5.3.2 Event and Tracking Data File Format
Figure 4 illustrates the first two events of the first half. Figures 5 shows a simplified example for two frames of player center of body tracking data. Note that we are only showing the mandatory CDF fields. These can be augmented with the optional fields (see Tables 9 and 10). If skeletal (limb) tracking is provided, it should follow the same structure as the tracking data JSON Lines format. Instead of a players $x$ and $y$ coordinates, the provider would provide coordinates for each landmark (e.g., $left\_hip\_x$, $right\_hip\_y$, $right\_hip\_x$) as specified by Table 5. The coordinate system should be the same used in the center of body tracking data, e.g., $x$, $y$, $z$ and are relative to the pitch. The field $is\_visible$ indicates if the part was actually measured, or inferred via an algorithm due to being hidden. Providers have flexibility over which body parts they track. This should be communicated beforehand.
@@ B.1 Optional Match Sheet Data
B.2 Optional Video Data
Table 8 lists optional video data.
Table 8: Description of the Optional Video Footage CDF fields.
Root Field Description Type recording/camera x x location of the camera in relation to the pitch center (m) Float recording/camera y y location of the camera in relation to the pitch center (m) Float recording/camera z z location of the camera in relation to the pitch center (m) Float
B.3 Optional Event Data
@@
B.4 Optional Tracking Data
@@
B.5 Optional Meta-Data